OpenStudy (anonymous):

can someone give me step by step how to solve this? Solve the equation: -2x-2+3x=6

OpenStudy (katrinakaif):

Combine like terms for x : So in this case -2x + 3x - 2 = 6 -2x + 3x = 1x 1x - 2 = 6 Bring the 2 over to the other side by adding it: 1x = 8 x = 8 when divided by 1 on both sides

OpenStudy (saifoo.khan):

-2x - 2 + 3x = 6 add like terms! x -2 = 6 x = 6 + 2 x = 8

OpenStudy (anonymous):

-4a-8\[\le\] 6a+22